The refractive index of the material of a double equiconvex lens is 1.5.if the radius of a curvature of the lens is r. Then its focal length is?

Grade:12

1 Answers

Gowtham
6 years ago
focal length is equal to its radius of curvature
from the lens maker formula
1/f=((n2-n1)/n1)(1/r1-1/r2)
here n1= refractive index of air
n2=refractive index of lens
r1 is positive and r2 is negative (from sign convention)
so 1/f=1/r
f=r